• Consider an economy described as follows: Y 5 C 1 I 1 G. Y 5 8,000. G 5 2,500. T 5 2,000. C 5 1000 1 2/3(Y2T ). I 5 1,200 2 100r. a. In this economy, compute private saving, public saving, and national saving. b. Find the equilibrium interest rate. c. Now suppose that G is reduced by 500. Compute private saving, public saving, and national saving. d. Find the new equilibrium interest rate


1
Expert's answer
2021-11-14T17:38:18-0500
"C=1,000+\frac{2}{3}Y-T"

"C=4,333"

"I=Y-C-G=1,167"

"1,167=1,200-100r"

"r=0.33"

If G=2,000


"I=1667"

"r=0.4"